3. Verizon is trading at $36. Put options with a strike price of $45 are priced at $10.50. What is the intrinsic value of the option, and what is the time value?  

4. Barclays plc is trading at £160 (Pound sterling). Put options with a strike price of £155 are priced at £5.5. What is the intrinsic value of the option, and what is the time value?
Answer #1

3

Intrinsic value = Max(strike price-current price,0) = Max(45-36,0) = Max(9,0) =9

Time value = Option price-intrinsic value = 10.5-9 = 1.5

4

Intrinsic value = Max(strike price-current price,0) = Max(155-160,0) = Max(-5,0) =0

Time value = Option price-intrinsic value = 5.5-0=5.5